OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)2 – 25? Identify the x–intercepts by graphing the quadratic equation. x–intercepts: (–2, 0) and (8, 0) y–intercept: (0, –16) x–intercepts: (2, 0) and (–8, 0) y–intercept: (0, –16) x–intercepts: (–2, 0) and (8, 0) y–intercept: (0, –25) x–intercepts: (2, 0) and (–8, 0) y–intercept: (0, –25)

OpenStudy (anonymous):

@RadEn @rebeccaskell94

OpenStudy (anonymous):

@nubeer

OpenStudy (anonymous):

first choice right ?

OpenStudy (raden):

first, x-intercept when y=0, so, (x-3)^2 - 25 = 0 (x-3)^2 - 5^2 = 0 (x-3+5)(x-3-5) = 0 (x+2)(x-8) = 0 for the zeroes, take each factor be zero : x+2=0 or x-8=0 x=-2 or x=8 thus, x-intercepts = {(-2,0),(8,0}

OpenStudy (raden):

then y-intercept when x=0 so, y = (x-3)^2 - 25 y = (0-3)^2 - 25 y = 9 - 25 y = -16 thus, y-intercept = (0,-16)